OpenStudy (anonymous):

Solve the system using substitution. x + y = 8 y = 3x

OpenStudy (anonymous):

look for x+y=8 for y we put 3x and it means that x +3x=8 4x=8 x=2

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

we know that y= 3x in another equation we only put not y but this 3x, that all that you need to do

OpenStudy (anonymous):

But x + 3x = 8 isn't an answer choice.

OpenStudy (brrandyn):

Okay, isolate a variable to one side of the equation. For instance:\[x + y = 8 \]Isolate:\[y=8-x\] Substitute y into the other equation! :) \[8 - x =3x\]Add x. \[8=4x\]Divide.\[2 = x\]

OpenStudy (brrandyn):

Is that easier to understand?

OpenStudy (anonymous):

I dont understand.

OpenStudy (okdutchman7):

Substitute 3x for y x + 3x = 8 Add the x's 4x = 8 Divide by 4 to get the x by itself x = 2 Plug 2 into the bottom equation for x y = 3(2) Multiply y = 6 Your answers: x = 2 y = 6

OpenStudy (anonymous):

you understood?

OpenStudy (brrandyn):

Look at the two equations you have. x + y = 8 y = 3x In the first equation x + y = 8, I moved x by subtracting it by both sides. So.. y = 8 - x <--- REPLACE Y IN THE OTHER EQUATION WITH THIS. Now SUBSTITUTE Y. LOOK AT YOUR SECOND EQUATION NOW. y = 3x ^ | 8 - x = 3x Add x. 8 = 4x Divide 4 by both sides. 2 = x

OpenStudy (okdutchman7):

@trenten 3 + 1

OpenStudy (anonymous):

lol yeah I understand now. Thanks

OpenStudy (brrandyn):

^ ^ .

OpenStudy (anonymous):

what? how was that hurtful? I said yours was the Best Response.. uh okay?

OpenStudy (anonymous):

how am I wrong?

OpenStudy (brrandyn):

Being sarcastic, do enjoy your evening.

OpenStudy (anonymous):

weird. okay.

OpenStudy (brrandyn):

If there's no other clarifying question, don't forget to close out. Though, it's at your discretion.